OpenStudy (anonymous):

Ben and John both drive 40km from home to work each day. One day ben said to john, "If you drive home at your usual speed, I will average 40kmph faster than you and arrive in 20 minutes less time." find John's speed. Please show full working, thanx

OpenStudy (anonymous):

Let Ben's speed be x and John's speed be y. x=y+40 Time taken for each to drive 40 km: Ben: 40/y = 40/(x+40) John: 40/x Time taken by Jonh is 20 minutes or 20/60 hours more than Ben Therefore, 40/x=40/(x+40) +20/60 40/x=40/(x+40)+1/3 40(1/x - 1/(x+40))=1/3 40*(x+40-x)/((x+40)*x)=1/3 4800=x^2+40x x^2+40x-4800=0 x=(-40+sqrt(1600+19200))/80 =(-40+sqrt(20800))/80 =(-40+40sqrt(13))/80 =(sqrt(13)-1)/2 y=x+40 = (sqrt(13)-1)/2 +40